If 30g of HCl is reacted with excess NaOH, and 10g of NaCl is produced, what is the theoretical yield of the experiment?
1
Expert's answer
2018-06-06T07:30:03-0400
HCl + NaOH → NaCl + H2O

n(HCl) = m/M = 30g/36.5g/mol = 0.82 mol

n(NaCl) = n(HCl) = 0.82 mol

M(NaCl) = 23 + 35.5 = 58.5 g/mol

m(NaCl) = n*M = 0.82mol*58.5g/mol = 47.97 g

47.97 g of NaCl with excess NaOH will theoretically produce 30 g of HCl.